OpenStudy (anonymous):

Find the polar coordinates of (-4,0) A. (0,0 degrees) B. (0,180 degrees) C. (4,180 degrees) D. None of the above

OpenStudy (klimenkov):

C.

OpenStudy (tyteen4a03):

@klimenkov Please don't give out answers without explaining. He would not be able to solve future questions like this without an explanation. To convert a Cartesian coordinate pair to a polar one (in the format \((r, \theta)\)), use this set of formula: r = \(\sqrt{x^2 + y^2}\) \(\theta = tan^{-1} (y/x)\).

OpenStudy (tyteen4a03):

Also, remember that the tangent value may be negative. In this case, apply the four quadrant rule on the final answer.

OpenStudy (klimenkov):

I don't think he wants to see how to solve this. Just want an answer.

OpenStudy (anonymous):

I am a she not a he :)

OpenStudy (tyteen4a03):

@klimenkov OpenStudy's Code of Conduct states that you should not give out answers without explanations. :)

OpenStudy (klimenkov):

I hope you will be satisfied now.